Solve for x by completing the square: 2x^2+8x-7=-2
Answer by ewatrrr(24785)

2x^2+8x-7=-2
2(x+2)^2 - 8 - 7 = -2
2(x+2)^2 = 13
(x+2)^2 = 13/2
x = -2 ± √13/2